OpenStudy (anonymous):

What is the x-intercept of the straight line having slope 3 and y-intercept 5

OpenStudy (anonymous):

since you know the a point (0, 5) and slope (m=3), you can use point slope form for the equation of the line in question: y - y1 = m(x - x1). then let y=0 in that equation and solve for x..

OpenStudy (anonymous):

a.)x-intercept = −5/3 b.) x-intercept =3/5 c.) x-intercept = −3/5 d.) x-intercept =5/3 e.) x-intercept = −5 f.) x-intercept = −3

OpenStudy (cwtan):

or in y=mx+c form y=3x+5 sub y=0 u will found x-intercept

OpenStudy (anonymous):

y=mx+b 0=3x+5 -5=3x x=-5/3 x intercept

OpenStudy (anonymous):

thank you !!!